OpenStudy (anonymous):

A company distributes free candies to all the students of x schools. Each school has (x + 1) classes. The number of students in each class is 3 more than the number of classes in each school. Each student is given 4 candies. Part A: Write an expression to show the total number of candies distributed by the company in x schools. (4 points) Part B: What would x(x + 1) represent? When simplified, what would be the degree and classification of this expression? (4 points) Part C: How can you calculate the total number of students in each school? (2 points)

OpenStudy (paki):

see... No. of schools = x No. of classes in each school = x+1 No. of students in each class = 3 more than number of classes in each school. Thus, no. of students in each class = (x+1) + 3 = x+4 No. of candies per student = 4... So the total number of candies is given 4 each to x+4 students in x+1 classes in x schools. So total number of candies = 4*(x+4)*(x+1)*x...

OpenStudy (paki):

For Part B) What would x(x+1) represent? When simplified, what would be the degree and classification of this expression? x is the number schools x+1 is the number of classes in each school So x*(x+1) would represent the total number of classes all taken together. When simplified x(x+1) = x2+1 which is a second degree polynomial... Part C may be solved by yourself for practice...

OpenStudy (anonymous):

I am having trouble solving part C

OpenStudy (anonymous):

At part A you know the number of candies being distributed, and since you know each student gets 4 candies, you can simply divide that by 4 and get the amount of students than right?

OpenStudy (paki):

its easy... try by making equation...